OpenStudy (anonymous):

the measures of each base angle of an isosceles triangle is 15 degrees less than the measure of the vertex angle. find the degree measure of the vertex angle. the measures of each base angle of an isosceles triangle is 15 degrees less than the measure of the vertex angle. find the degree measure of the vertex angle. @Mathematics

OpenStudy (anonymous):

x = vertex angle base angle = (180-x)/ 2 x - 15 = (180-x)/2 2x - 30 = 180-x 3x = 150 x = 50